# Using $k$-medoids for distributed approximate similarity search with arbitrary distances

### v1

@@ 3 Methodology
This section describes the methodology developed for data indexation using arbitrary distance functions. First, it illustrates the GMASK algorithm as a general approximate indexing method accepting any distance function. Hereafter, it delves into the characteristics of Minkowski distances, emphasising the inherent limitations in $k$-means that justify the implementation of GMASK. Consequently, we justify adopting $k$-medoids, an alternative clustering method that leverages the opportunities provided by Minkowski distances to overcome previous restrictions derived from using the Euclidean distance. The design of experimental tests is described in Subsections “Performance evaluation”, “Benchmark algorithms for similarity search”, and “Datasets”, including metrics to assess performance, state-of-the-art algorithms for comparison, and the datasets selected to gauge effectiveness.
3.1 GMASK
As mentioned above, there is a close connection between machine learning and similarity search indexing. Firstly, the foundations for designing data access methods rely on familiar concepts in machine learning classification, such as dissimilarity spaces and distance functions. Secondly, clustering methods like $k$-means or $k$-medoids are used in many of these algorithms to create indexing structures for similarity search and improve search performance. Besides, they are also affected by the same limitations, such as the Curse of Dimensionality and the challenge of identifying the actual intrinsic dimensionality governing some problems to make them solvable.
The typical choice in partitioning clustering methods is to find a low value of $k$ that generates different groups in the dataset. GMASK follows an unconventional approach for clustering, generating a high number of $k$ representative points for each level. The only requirement in GMASK is using a clustering method that generates Voronoi regions and returns a representative point for each of them. In this implementation, GMASK uses the $k$-medoids clustering method, inducing a Voronoi tessellation and returning an element representing each region. Thus, generating a large number of representative points, GMASK summarises the underlying data points with finer detail. In the same way, this indexing method applies to any dissimilarity space, including the particular case of metric spaces or vector spaces. Thanks to the $k$-medoids adaptation in GMASK, this algorithm presents an important improvement over MASK: it is able to work with different distance functions. This modification brings GMASK the capability of adaptation to the specific characteristics of any dataset. In the “Experimental results” section, we show that this unusual application of clustering has attractive properties and yields superior performance, even with high-dimensional datasets residing in feature spaces that are typically sparse.
At this point, it is essential to note that the goal of GMASK is not to cluster data but to use the partitioning clustering method to construct a multilevel index to find elements. With more powerful computational infrastructures and large memory and storage capacity becoming prevalent, the advantage of this method lies in that previous restrictions on the number of clusters at each level become less relevant. In this new scenario, we can exploit the effect of introducing a high number of prototype points into the dataset, leveraging the capabilities of the computing infrastructure.
Figure 2 illustrates the design principles of GMASK. At the lowest level, we have the data points in $\mathcal{X}$ to be indexed. The blue boxes in the diagram represent different data partitions or groups that could be stored on various computing nodes. The novelty lies in creating the multilevel structure of $k$ medoids, following a bottom-up approach. The first level of medoids (identified as $n_{1}$ in Figure 2) summarises the data points at the lowest level so that points assigned to each cluster are represented by its prototype (medoid). The second level (denoted $n_{2}$) corresponds to medoids representing the prototypes computed at the first level, $n_{1}$. The same strategy is repeated, creating additional prototype layers that summarise the immediately lower group of medoids. The procedure stops when the number of medoids at the top level of this index structure becomes manageable.
Figure 2: Conceptual representation of the updated multilevel structure approach created by GMASK.
The process of constructing this multilevel index structure depends on two initial parameters:
@@ 3.1 GMASK
  • The group length quantifies the number of elements in each data partition at the lowest level, configured according to the available computing power (and the total number of nodes in distributed computing systems). For the baseline, we assume that all partitions are of the same size, although this condition is not strictly necessary.
@@
  • The number of medoids calculated for each group. At this point, the question of selecting the most optimal number of medoids has not yet been addressed, but it is not very important. Contrary to the standard procedure, our approach suggests that, for information indexing, one should use as many medoids as the computational infrastructure can afford.
The clustering at the first level reduces the initial number of data points by a particular proportion, the data summarisation ratio, which is determined by the ratio between the group length and the number of medoids. This ratio also determines the number of layers of the multilevel index and the stop condition when the number of medoids at the top layer becomes lower than the group length.
Each medoid is a prototypical point representing all the elements assigned to its cluster. The number of points assigned to each cluster may vary depending on their distance from the corresponding medoid. For this reason, regions with higher data density will also tend to receive more medoids. By generating a high number of medoids, we can capture finer details of the underlying data, allowing us to index not only high-density regions but also outliers and clusters of elements distinct from the main concentrations of points. Algorithm 1 describes the construction of this multilevel index structure.
Algorithm 1 Multilevel index construction in GMASK
1:function Clustering($distFunc,nPrototypes,points$)
2:     $\triangleright$ A clustering method returning Voronoi regions
3:     $\triangleright$ and a prototype point representing each region
4:end function
5:
6:procedure GMASK($data,lengthGroup,nPrototypes$)
7:     $ngroups\leftarrow lengthData/lengthGroup$;
@@
8:     $vector\leftarrow$ split($data,nGroups$); $\triangleright$ Divide dataset into nGroups
9:     $\triangleright$ Initialise lists $layerPoints$, $layerLabels$, and variable $idLayer$
10:     while $ngroups\geq 1$ do
@@
11:         $\triangleright$ Initialise groupsPoints, groupsLabels and points
12:         for $idGroup=1,nGroups$ do
13:              $points\leftarrow vector[idGroup]$;
14:              $groupsPoints.add($clustering($distFunc,nPrototypes,points$));
15:              $groupsLabels.add(clusters.labels)$;
16:         end for
17:         $layerPoints.add(groupsPoints)$;
18:         $layerLabels.add(groupsLabels)$;
19:         $ngroups\leftarrow length(groupsPoints)/lengthGroup$;
20:         if $ngroups\geq 1$ then
21:              $vector\leftarrow$ split($layerPoints,nGroups$);
22:         end if
23:         $idLayer\mathrel{+}=1$;
24:     end while
25:     $layers\leftarrow idLayer-1$;
26:end procedure
Once built according to the rules described above, the index is ready to accept point queries, which systematically search for all points in the original dataset, seeking a perfect match. To solve any query using this multilevel index, we start at the top level of the medoids hierarchy, steering the search according to the distance between the medoids at each level and the target query object $q\in\mathcal{X}$.
3.2 Dissimilarity measures
A dissimilarity space [26] defines the rules for measuring the distance between elements represented in a feature space. As discussed in Section “Similarity search”, similarity captures the notion of closeness between two objects. Euclidean vector spaces are widely applied in previous research, as they reflect and intuitive interpretation of a space: vectors are elements that can be added, multiplied by a scalar or projected onto each other.
Being $\mathbb{X}$ a set of elements, the function
$\delta:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ (1)
returns a real number $\delta(a,b)$ that shall be used to quantify the difference between two or more elements contained on this set, denoting a measure of distance.
Distance functions have the following properties:
  • $(p1)\quad\forall a,b\in\mathbb{X},\delta(a,b)\geq 0$ positiveness,
  • $(p2)\quad\forall a,b\in\mathbb{X},\delta(a,b)=\delta(b,a)$ symmetry,
  • $(p3)\quad\forall a\in\mathbb{X},\delta(a,a)=0$ reflexivity,
  • and in most cases
  • $(p4)\quad\forall a,b\in\mathbb{X},a\neq b\to\delta(a,b)>0$ strict positiveness.
If it also satisfies
  • $(p5)\quad\forall a,b,c\in\mathbb{X},\delta(a,b)\leq\delta(a,c)+\delta(c,b)$, commonly known as triangle inequality,
$\delta$ is indeed a metric, and the pair $(\mathbb{X},\delta)$ is called a metric space [6]. If the distance does not satisfy the strict positiveness property $(p4)$ then the space is called a pseudo-metric space, as well as if it does not hold the symmetry property $(p2)$, we may have a quasi-metric space. Nevertheless, there are some primitive spaces where these metric constraints are not necessarily satisfied. We also use the word ’dissimilarity’ when referring to this difference or distance between elements [18].
Normed and inner product spaces are special cases of metric spaces, where a metric is defined either by a norm or an inner product. An Euclidean space is a specific case of a Hilbert space. In turn, a Hilbert space is an inner product space and a particular case of a Banach space, an example of normed spaces in which a norm metric $\ell norm$ has been defined [5].
3.3 Minkowski distance
@@
Minkowski distances or $\ell_{p}norms$ are distance measures commonly used to compute the dissimilarity of objects described by numerical attributes. However, dissimilarity can also be quantified using non-numerical attributes through other alternatives, such as the edit distance widely used for text attributes.
@@
3.4 Limitations of k-means and the Euclidean distance
@@
In unsupervised machine learning with $k$-means, the user must determine the appropriate number of clusters for a given problem, and a centroid represents each cluster. Carefully choosing the number of centroids is crucial because too many centroids would not effectively summarise the underlying data, and too few centroids would result in grouping unrelated data points. However, the locations provided to these initial centroids can greatly impact the final outcome, making it a sensitive aspect.
@@
To address this limitation, algorithms such as $k$-means++ [27] propose an approach where the initial random centroids are spread more evenly, leading to improved results. Nevertheless, although several more powerful $k$-means implementations have tried to reduce the performance impact resulting from sub-optimal initial centroid placement, only a few have explored the potential for achieving better solutions by choosing a more appropriate distance.
Due to the nature of the $k$-means algorithm, the choice of distance function can also exert a significant impact on $k$-means outcomes, as it determines how similarity between data points is calculated and produces clusters with different shapes and separations between each other. As it has been raised in Section “Background and related work”, despite the widespread use of Euclidean distance in indexing structures and algorithms, relying on it as the default option when using the $k$-means algorithm can lead to poor results.
@@
Previous research [22] suggests that inherent characteristics of high-dimensional spaces, such as high dispersion, outliers or concentration of points around specific values, hinder the discriminatory power of the $\ell_{p}norm$ for higher values of $p$. This situation would provide undue importance to the distant (sparse/noisy) dimensions and poorly discriminate between further and nearer neighbours. Therefore, for a given high-dimensional problem, it may be preferable to use Minkowski distances with lower values of $p$, so the $\ell_{1}$ distance metric (Manhattan distance) would be the best choice for high-dimensional applications, followed by the Euclidean distance ($\ell_{2}$), then the $\ell_{3}$ metric, up to the Chebyshev distance ($\ell_{\infty}$).
@@
Previous algorithms for approximate similarity search, such as MASK [2], have relied on the Euclidean distance to calculate relative dissimilarities between points. This strategy follows the conventional inclination to associate distance between points with the Euclidean metric, which aligns with the distance concept commonly applied in geospatial problems. However, the potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a broader range of dissimilarity measures [5], justify our efforts to explore alternative Minkowski distances for data indexing. This endeavour aims to foster a deeper understanding of the impact of using alternative distance metrics with high-dimensional data.
@@
Nevertheless, while it may seem theoretically plausible to employ $k$-means or its enhanced version, $k$-means++, with several distance metrics, the practical implementation reveals that the algorithm’s internal design is intertwined with the Euclidean distance. This alignment relies on the algorithm’s iterative optimisation process when it searches for centroids that minimise the total squared distances of points within their respective clusters[3]. The sum of the Squared Error (SSE), calculated as the sum of squared Euclidean distances between each data point and the centroid of its assigned cluster, underscores the algorithm’s inherent reliance on Euclidean geometry for clustering optimisation [28].
As we have already demonstrated, assuming that widely adopted mechanisms are the most suitable or the only viable option can lead to self-imposed limitations that generate sub-optimal outcomes. Instead, our objective is to steer clear of the misconception that certain implementations are the best or the only conceivable choice simply because they are the most common choices.
@@
At this juncture, GMASK can integrate an alternative clustering algorithm that leverages the opportunities provided by diverse distance metrics for data indexing. The $k$-means algorithm defines clusters based on centroids minimising the sum of squared errors in distances from all other points in the same cluster. In contrast, alternative methods like $k$-medoids calculate clusters around the most representative data element, which is determined by minimising the sum of pairwise dissimilarities among data points the same cluster.
3.5 k-medoids as a generalisation of k-means
Several generalisations of the $k$-means algorithm have been proposed to broaden the scope of applicable data types [29], reflecting the diversity found in nature prototype points, such as subsets of objects, hyperplanes, etc.
@@
The first paper proposes a general criterion and its generalised $k$-means method is [30]. This approach starts from a $n$ x $n$ dissimilarity matrix $(d_{\ell}t)$ for $n$ factories $O=1,...,n$ in an industrial network where $d_{\ell}t$ are the minimum road transportation costs between $\ell$ and $t$. The goal is to partition the set of factories into $k$ classes $(C_{1},...,C_{k})$ and to find a selection $Z=(z_{1},...,z_{k})$ of $k$ factories as ’supply points’ such that, when supplying all factories of the class $C_{i}$ from the supply point $z_{i}\in O$, the overall transport costs are minimised, where $d(\ell,z_{i})=d_{\ell},z_{i}$ means the dissimilarity between the factory (object) $\ell$ and the factory (supply point) $z_{i}\in O$. So the family $P$ of admissible ’prototypes’ consists of all singletons from $O$ and the class prototypes, meaning ’most cheapest supply point’ in $C_{i}$, are determined by:
@@
Many authors, including, among others, Diday [31, 32, 33], Bock [34] or Govaert [35], have followed the generalised clustering approach in various settings and numerous variations, obtaining a plethora of generalised $k$-means algorithms.
Kaufmann and Rousseeuw [36] first propose the popular Partitioning Around Medoids (PAM) algorithm, also referred to as $k$-medoids. In $k$-medoids, data is modelled by using $k$ representative objects $m_{i}$ called medoids, that serve as prototypes for the clusters (formerly called classes) which are obtained by using the absolute error criterion (’total deviation’, TD). Since the TD is defined as the sum of dissimilarities of each point $x_{j}\in C_{i}$ to the medoid $m_{i}$ of its cluster, the medoid of a cluster $C_{i}$ is in fact the object with the smallest sum of dissimilarities (or equivalently, the smallest average) to all other objects in the cluster.
@@
This method is suitable for non-Euclidean distances and a wide range of data since it does not force the dissimilarity measure to adhere strictly to metric properties, and only requires a proximity measure between any pair of objects. In addition, while $k$-means defines a prototype in terms of a centroid, which is usually the mean of a group of points that may not correspond to an element in the dataset, $k$-medoids defines a prototype in terms of a medoid, an actual data point identified as the most representative one for a group of points.
Regarding $k$-medoid implementations, Program PAM [36] consists of two algorithms, BUILD to choose an initial clustering, and SWAP to improve the clustering towards a local optimum, since finding the global optimum of the $k$-medoids problem is, unfortunately, NP-hard. The algorithms require a dissimilarity matrix, which requires $O(n^{2})$ memory and at least $O(n^{2}d)$ time to be computed. On top of this, the process of updating medoids involves evaluating the cost of swapping non-medoid points to ensure that the sum of pairwise dissimilarities within clusters (rather than the sum of squared Euclidean distances) is minimised, making this method more robust to noise and outliers compared to the alternatives based on $k$-means. Recently, FasterPAM [37] introduced a speedup for larger $k$, achieving an $O(k)$-fold acceleration in the second SWAP phase of the algorithm, but still providing the same results as the original PAM algorithm.
Originally FasterPAM was implemented in Java and published within the open-source library ELKI [38]. However, to integrate it with GMASK implementation, developed entirely in Python, we will employ the fast $k$-medoids Python implementation [39] version provided by the $kmedoids$ package  [^1] https://pypi.org/project/kmedoids. FasterPAM let us upgrade our method by making it compatible with any distance function and a wider range of problems.
Experimental results verify the applicability of GMASK with real datasets, improving the performance of alternative algorithms and providing better results for approximate similarity search queries.

### v2

@@ 3 Methodology
This section describes a new data indexing framework called GDASC. First, it illustrates the properties and structure of GDASC. Then, it delves into the characteristics of different distance metrics, emphasising the inherent limitations in $k$-means. Aside from this traditional clustering algorithm, GDASC accepts other clustering algorithms such as $k$-medoids, which can exploit the properties of non-Euclidean or even non-metric distance functions for specific datasets. The design of experimental tests is described in Subsections 3.4, 3.5, and 3.6, including the metrics used to assess performance, state-of-the-art algorithms for comparison, and the datasets selected to gauge their effectiveness.
3.1 GDASC
GDASC is a versatile framework for data indexing as it can adopt different clustering algorithms and distance functions. Figure 2 depicts the three main blocks of the framework: Dataset Description (DD), Multilevel Structure Algorithm (MSA) and the Neighbours Search Algorithm (NSA). Besides the Dataset Description, GDASC relies on two algorithms, one for building a multilevel index structure (MSA) and another one for searching neighbours (NSA), described in Algorithm 1 and Algorithm 2 respectively. The proposed approach for building the multilevel structure, which is used in the search process, is prepared to receive any clustering algorithm that generates representatives summarising the underlying dataset. This is an important feature of GDASC since it is possible to adapt the framework to the particular characteristics of the dataset.
Figure 2: GDASC framework is composed of three main parts.
Given that one of the main contributions of GDASC is its applicability to real-world datasets with minimal or no preprocessing, the DD module performs an initial descriptive analysis of the dataset before conducting any experiments. Its goal is to determine adequate values for the input parameters of MSA and NSA. Firstly, by examining the number of elements and dimensions of the dataset, we can determine if the problem is described in a high-dimensional space or is computationally unfeasible. This assessment may lead us to consider distributing the dataset across multiple computational nodes. Next, for each dimension of the dataset, we will calculate the range of values to determine its sparsity and to gain an initial understanding of the distribution of points. We will also compute measures of central tendency (mean and median), dispersion (standard deviation), and distribution shape (kurtosis and skewness). Finally, for each distance metric to be applied for similarity search, we calculate the maximum and minimum distances between two points in the dataset, the average distance between points, and the distribution of distance quantiles. These calculations provide a deeper understanding of the dataset and help determine the values of initial parameters that depend on dataset features, such as the radius.
The MSA module uses Algorithm 1 to build a multilevel index structure following a bottom-up approach and using a cluster algorithm with an initial number of prototypes as high as possible, depending on the computational capacity. In the use of clustering algorithms, is very common to find a low number of centroids that generates different groups in the dataset. However, MSA follows an rather unusual approach for clustering, generating instead a high number of representative points for each level (prototypes). This way, the resulting indexing structure can summarise the underlying data points with finer detail.
The structure generated by the MSA module depends on several initial parameters:
@@ 3.1 GDASC
  • The group length ($gl$) quantifies the number of elements in each data partition at the lowest level, configured according to the available computing power (and the total number of nodes in distributed computing systems).
@@
  • The number of prototypes ($nPrototypes$) calculated for each group. Contrary to the standard procedure, our approach suggests that, for information indexing, one should use as many prototypes as the computational infrastructure can afford. Empirically we have concluded that the ratio between the group length and the number of prototypes have to be 2:1. This ratio also determines the number of layers of the multilevel index.
  • The distance function ($distFunc$) applied by the clustering algorithm. This is a critical decision because not all distance functions work properly with any dataset.
  • The clustering algorithm ($clusterAlg$). Clustering methods like $k$-means or $k$-medoids are widely used in algorithms that create indexing structures for similarity search. However, MSA is able to work with any clustering algorithm. In this paper, we focus in the use of $k$-medoids and consequently employ various distance functions.
Algorithm 1 Multilevel Structure Algorithm in GDASC
1:procedure MSA($data,gl,nPrototypes,distFunc,clusterAlg$)
@@
2:     $vector,nGroups\leftarrow$ split($data,gl$); $\triangleright$ Divide dataset into groups of $gl$
3:     $\triangleright$ Initialise multilevel structure: $levelPoints$, $levelLabels$, and variable $idLevel$
4:     while $nGroups\geq 1$ do
@@
5:         $\triangleright$ Initialise auxiliary structures for each level: $groupsPoints$, $groupsLabels$ and $points$
6:         for $idGroup=1,nGroups$ do
7:              $points\leftarrow vector[idGroup]$;
8:              $groupsPoints.add($clustering($clusterAlg,distFunc,nPrototypes,points$));
9:              $groupsLabels.add(clusters.labels)$;
10:         end for
11:         $levelPoints.add(groupsPoints)$;
12:         $levelLabels.add(groupsLabels)$;
13:         $vector,nGroups\leftarrow$ split($levelPoints,gl$);
14:         $idLevel\mathrel{+}=1$;
15:     end while
16:     $nLevels\leftarrow idLevel-1$;
17:end procedure
Figure 3 illustrates the schema implemented by MSA. At the lowest level, we have the data points in $\mathcal{X}$ to be indexed. They are divided into several groups according to the $gl$ parameter. Once these groups are formed, the points in the next level are the representative points of the previous level. That is, following a bottom-up approach, the first level of prototypes (identified as $n_{1}$ $prototypes$ in Figure 3) summarises the data points at the lowest level so that points assigned to each cluster are represented by its prototype. The second level (denoted $n_{2}$ $prototypes$) corresponds to prototypes representing the prototypes computed at the first level, $n_{1}$ $prototypes$. The same strategy is repeated, creating additional prototype layers that summarise the immediately lower group of prototypes. The procedure stops when the number of prototypes at the top level of this index structure becomes manageable.
It can be remarkable that the groups formed at the lower level or the cluster formed at any level can have distinct size. When the number of points in a group or a cluster is less or equal than the number of prototypes, all the points are become prototypes. In the next level, the groups or prototypes are gather in groups with, at least, as points as $gl$. At level 0, the last three groups have less points than $gl$. Thus, in level 1, the prototypes of these groups are arranged in one group and five prototypes are selected for level 2.
Figure 3: Conceptual representation of the updated multilevel structure approach created by GDASC.
The points assigned to each cluster may vary depending on the distance function applied and its value. For this reason, regions with higher data density will also tend to receive more prototypes. By generating a high number of prototypes, we can capture finer details of the underlying data, allowing us to index not only high-density regions but also outliers and clusters of elements distinct from the main concentrations of points.
The NSA module uses the multilevel index structure generated by MSA to find the $k$ nearest neighbours by applying Algorithm 2. For a given point query $(q)$, a number of neighbours $(k)$ and a radius, this algorithm goes from the top level of the multilevel structure to the bottom one, testing the distance between prototypes and the query point. In each level, the algorithm computes the distance between the point query and the prototypes in that level. Only the prototypes whose distance is less than radius will be consider in the next level. Therefore, the search way goes throw prototypes that satisfy this filtering criterion.
Algorithm 2 Neighbours Search Algorithm in GDASC
procedure NSA($levelPoints,levelLabels,nLevels,distFunc,q,k,radius$)
2:     $prototypes\leftarrow levelPoints[nLevels]$;
     $D\leftarrow$ distance($distFunc,q,prototypes$);
4:     $idCandidates\leftarrow D<radius$;
     for $level=nLevels-1,1$ do
6:         $prototypes\leftarrow levelPoints[levels][idCandidates]$;
         $D\leftarrow$ distance($distFunc,q,prototypes$);
8:         $idCandidates\leftarrow D<radius$;
         $level\mathrel{1}=1$;
10:     end for
     $candidates\leftarrow levelPoints[0][idCandidates]$;
12:     $D\leftarrow$ distance($distFunc,q,candidates$);
     $candidates\leftarrow$ sort($candidates,D$);
14:     $neighbours\leftarrow candidates[0:k]$;
end procedure
Figure 4 represents the search procedure of $k=2$ neighbours using a multilevel structure with 4 levels. In a previous step, NSA computes the distance between the point query $q$ and the prototypes at level 3, $d(C_{3-1},q),d(C_{3-2},q),d(C_{3-3},q)$. As only $d(C_{3-1},q)$ and $d(C_{3-3},q)$ are less than radius, only the points in $C_{3-1}$ and $C_{3-2}$ clusters are considered. Thus, the algorithm calculates the distance between $q$ and all the points in $C_{3-1}$ and $C_{3-3}$ clusters and only selects the points for which this value is less than radius. Paying attention to $C_{3-1}$, this cluster contains three points and only the distance for two of them are less than radius. Therefore, the arrows between $C_{3-1}$ and these points are represented by a solid pink line, while the other arrow is represented by a dotted grey line. In the next level, the algorithm only analyses the distances for the points in $C_{2-1}$ and $C_{2-3}$ clusters. When the bottom level is reached, the algorithm has selected several points as potential candidates. They are filtered according to the distance between them and $q$ selecting the $k=2$ points with less distance value.
Figure 4: Search algorithm example.
3.2 Distance metrics
Distance metrics provide a quantitative means of assessing the dissimilarity between objects represented in a feature space. In this paper we focus on two specific distance metrics: Minkowski distance, which generalises multiple distance measures, and cosine distance, which assesses the dissimilarity between vectors based on their angle in a high-dimensional space.
Minkowski distance
@@ 3.2 Distance metrics
The Minkowski distance or $\ell_{p}norm$ conforms a set of distance measures commonly used to compute the dissimilarity of objects described by numerical attributes. However, dissimilarity can also be quantified using non-numerical attributes through other alternatives, such as the edit distance, widely used for text attributes.
@@
Cosine distance
The cosine distance is a measure of dissimilarity between two non-zero vectors in an inner product space, which calculates the cosine of the angle between them. It is a handy distance in high-dimensional spaces and is widely applied in many fields such as text mining, information retrieval, and machine learning, especially in document similarity and clustering [23]. Unlike the Euclidean distance that considers the magnitude of the vectors, cosine distance focuses on their orientation, making it less sensitive to differences in vector length and more robust to variations in data scaling [46]. Cosine distance has two main differences from Euclidean distance. Firstly, it is scale-invariant because it measures the angle between two vectors rather than their magnitude. Secondly, it assigns higher weight to dimensions with higher values, whereas Euclidean distance treats all dimensions equally.
The cosine distance measures the dissimilarity between two non-zero vectors and can be formulated as one minus the cosine similarity between them. This property makes it particularly effective in handling sparse data and datasets with many features, where traditional distance metrics often struggle due to the Curse of Dimensionality. Mathematically, for two vectors A and B, the cosine distance is defined as:
$Distance_{A,B}=1-cos(\theta)=1-\frac{A\cdot B}{\|A\|\|B\|}=1-\frac{\sum_{i=1}^{n}A_{i}B_{i}}{\sqrt{\sum_{i=1}^{n}A_{i}^{2}}\sqrt{\sum_{i=1}^{n}B_{i}^{2}}},$ (2)
where $\theta$ is the angle between A and B in the n-dimensional space, and $A_{i}$ and $B_{i}$ are the $i$-th components of vectors A and B, respectively. Thus, the cosine distance ranges from 0 to 2, where 0 means that the vectors are identical and 2 means that they are diametrically opposed to each other.
3.3 Clustering algorithms
@@
There are many clustering algorithms for unsupervised machine learning and one of the most common options is $k$-means. When applying $k$-means the user must usually determine the appropriate number of clusters for a given problem, and the algorithm obtains a centroid representing each cluster. Choosing the number of clusters is crucial because too many centroids would not effectively summarise the underlying data and too few centroids would result in grouping unrelated data points. Besides, the locations assigned to these initial centroids can greatly impact the final outcome, making it a sensitive aspect.
@@
To elude this limitation, algorithms such as $k$-means++ [50] propose an approach where the initial random centroids are spread more evenly, leading to improved results. Nevertheless, although several more powerful $k$-means implementations have tried to reduce the performance impact resulting from sub-optimal centroid initialisation, only a few have explored the potential for achieving better solutions by choosing a more appropriate distance.
@@
The choice of distance function to be used with $k$-means, can also exert a significant impact since it determines how similarity between data points is calculated and produces clusters with different shapes and separations. Previous research [46] suggests that inherent characteristics of high-dimensional spaces, such as high dispersion, outliers or high concentration of points around specific values, hinder the discriminatory power of the $\ell_{p}norm$ for higher values of $p$. This situation would provide undue importance to the distant (sparse/noisy) dimensions and poorly discriminate between further and nearer neighbours, making it preferable to use Minkowski distances with lower values of $p$. For instance, the $\ell_{1}$ distance metric (Manhattan distance) would be a better choice for high-dimensional applications, followed by the Euclidean distance ($\ell_{2}$), then the $\ell_{3}$ metric, and so for, up to the Chebyshev distance ($\ell_{\infty}$).
@@
As we have indicated in Section 2, despite the widespread use of indexing structures and algorithms based on the Euclidean distance [51], as a default option when using the $k$-means algorithm, such a choice can lead to poor results. This strategy follows the conventional inclination to associate distance between points with the Euclidean metric, which aligns with the distance concept commonly applied in geospatial problems. However, the potential for achieving improved results, coupled with the lack of prior empirical studies validating this proposal, and the need for learning paradigms applicable to a broader range of dissimilarity measures [9], justify our efforts to explore non-Euclidean distances for data indexing. This endeavour fosters a deeper understanding of the impact of using alternative distance functions with high-dimensional data.
@@
Nevertheless, while it may seem theoretically plausible to employ $k$-means or its enhanced version, $k$-means++, with several distance metrics, the practical implementation reveals that the algorithm’s internal design is intertwined with the Euclidean distance. This alignment relies on the algorithm’s iterative optimisation process, when searching for centroids that minimise the total squared distances of points within their respective clusters[52]. The sum of the Squared Error (SSE), calculated as the sum of squared Euclidean distances between each data point and the centroid of its assigned cluster (so-called Within-Cluster Sum of Squares or WCSS), underscores the algorithm’s inherent reliance on Euclidean geometry for clustering optimisation [53].
@@
As we have already demonstrated, assuming that widely adopted mechanisms are the most suitable or the only viable option can lead to sub-optimal outcomes. Instead, our goal is to steer clear of the misconception that certain implementations are the best or the only conceivable choice because they are the most common. Thus, GDASC can integrate alternative clustering algorithms to create the indexing structure. While $k$-means defines clusters based on centroids minimising the sum of squared errors in distances from all other points in the same cluster, alternative methods like $k$-medoids calculate clusters around the most representative data element, which is determined by minimising the sum of pairwise dissimilarities among data points in the same cluster.
@@
The first paper proposing a general criterion and a generalised $k$-means method is [54]. This approach starts from a $n$ x $n$ dissimilarity matrix $(d_{\ell}t)$ for $n$ factories $O=1,...,n$ in an industrial network where $d_{\ell}t$ are the minimum road transportation costs between $\ell$ and $t$. The goal is to partition the set of factories into $k$ classes $(C_{1},...,C_{k})$ and to find a selection $Z=(z_{1},...,z_{k})$ of $k$ factories as ’supply points’ such that, when supplying all factories of the class $C_{i}$ from the supply point $z_{i}\in O$, the overall transport costs are minimised, where $d(\ell,z_{i})=d_{\ell},z_{i}$ means the dissimilarity between the factory (object) $\ell$ and the factory (supply point) $z_{i}\in O$. So the family $P$ of admissible ’prototypes’ consists of all singletons from $O$ and the class prototypes, meaning ’most cheapest supply point’ in $C_{i}$, are determined by:
@@
Many authors, including Diday [1971, 1973, 1974], Bock [1970] or Govaert [59], have followed the generalised clustering approach in various settings and numerous variations, obtaining a plethora of generalised $k$-means algorithms.
Kaufmann and Rousseeuw [1987] first proposed the popular Partitioning Around Medoids (PAM) algorithm, also referred to as $k$-medoids. Data is modelled using $k$ representative objects $m_{i}$ (called medoids), that serve as prototypes for the clusters (formerly called classes). These prototypes are obtained using the absolute error criterion (’total deviation’, TD). Since the TD is defined as the sum of dissimilarities of each point $x_{j}\in C_{i}$ to the medoid $m_{i}$ of its cluster, the medoid of a cluster $C_{i}$ is in fact the object with the smallest sum of dissimilarities.
@@
This method is suitable for non-Euclidean distances and a wide range of data since it does not force the dissimilarity measure to adhere strictly to metric properties, and only requires a proximity measure between any pair of objects. While $k$-means defines a prototype in terms of a centroid, $k$-medoids defines a prototype in terms of a medoid, an actual data point identified as the most representative one for a group of points.
As for $k$-medoids implementations, Program PAM [60] consists of two algorithms: BUILD to choose an initial clustering and SWAP to improve the clustering towards a local optimum, since finding the global optimum of the $k$-medoids problem is NP-hard. The algorithms need a dissimilarity matrix, which requires $O(n^{2})$ memory and at least $O(n^{2}d)$ time to be computed. On top of this, the process of updating medoids involves evaluating the cost of swapping non-medoid points to ensure that the sum of pairwise dissimilarities within clusters (rather than the sum of squared Euclidean distances) is minimised, making this method more robust to noise and outliers compared to the alternatives based on $k$-means. Recently, FasterPAM [61] introduced a speedup for larger $k$, achieving an $O(k)$-fold acceleration in the second SWAP phase of the algorithm, but still providing the same results as the original PAM algorithm.
Originally, FasterPAM was implemented in Java and published within the open-source library ELKI [62]. However, to integrate it with GDASC implementation, developed entirely in Python, we will employ the fast $k$-medoids Python implementation [63] version provided by the $kmedoids$ package  [^1] https://pypi.org/project/kmedoids. FasterPAM let us upgrade our method by making it compatible with any distance function and a wider range of problems.